# How Far Can Transformers Reason? The Locality Barrier and Inductive Scratchpad

### v1

@@ Abstract
Can Transformers predict new syllogisms by composing established ones? More generally, what type of targets can be learned by such models from scratch? Recent works show that Transformers can be Turing-complete in terms of expressivity, but this does not address the learnability objective. This paper puts forward the notion of distribution locality to capture when weak learning is efficiently achievable by regular Transformers, where the locality measures the least number of tokens required in addition to the tokens histogram to correlate nontrivially with the target. As shown experimentally and theoretically under additional assumptions, distributions with high locality cannot be learned efficiently. In particular, syllogisms cannot be composed on long chains. Furthermore, we show that (i) an agnostic scratchpad cannot help to break the locality barrier, (ii) an educated scratchpad can help if it breaks the locality at each step, (iii) a notion of ‘inductive scratchpad’ can both break the locality and improve the out-of-distribution generalization, e.g., generalizing to almost double input size for some arithmetic tasks.
@@ 1.2 Hardness of long compositions
Consider the previous syllogism composition task where implications are drawn on a graph with 24 edges drawn randomly over 24 vertices. Picking vertices at distances 1 to 4 for the connected case and picking disconnected vertices uniformly at random lets a Transformer achieve a test accuracy of more than 80% after about 2K iterations. However, does this mean that the model has learned to compose syllogisms, or has it found shortcuts, e.g., based on node degrees, to guess the implications often enough? In Appendix B.1, we provide empirical evidence supporting the latter. Motivated by this issue and to preclude spurious correlations, we consider the following distribution.
@@
(b) Nb. of iterations (y-axis) with $512$ fresh samples to learn the cycle task of size $n$ (x-axis) using GPT2-style models with $10M,25M,85M$ parameters.
@@
A challenge for the cycle task is that there is no clear ‘low-complexity pattern’ in the input representation that indicates whether there are 1 or 2 cycles. No simple statistics based on degrees, edge counts, or finite motif counts that can tell if the vertices are connected or not. One has to consider at least $n$ edges in order to get any correlation with the presence of a path. In other words, the task requires a ‘global reasoning’ involving a ‘large’ number of input tokens and this seems hard for Transformers.
@@ 1.3 Hardness of global reasoning
As discussed previously, the cycle task appears to be challenging for Transformers as it requires some global reasoning. Other tasks such as subset parities exhibit the same challenge. However the latter can be proved to be not efficiently learnable by various regular neural networks and noisy gradient descent, as one can get explicitly a class of functions (through orbit arguments [12, 13]) that has large statistical dimension [14] or low cross-predictability [12, 15] (see Appendix A.2). For the cycle task, we have a single distribution, and it is unclear how to use the invariances of Transformers to get arguments as in [12, 13], as the input distribution is not invariant under the symmetries of the model. We thus would like to develop a more general complexity measure that unifies why such tasks are hard for Transformer-like models and that formalizes the notion of ‘local reasoning barrier’ when models are trained from scratch. We also would like to understand how the scratchpad methodologies that have proved helpful in various settings (see Section 3) can help here. This raises the questions:
(1) How can we formalize the ‘local reasoning barrier’ in general terms?
(2) Can we break the ‘local reasoning barrier’ with scratchpad methodologies?
@@ 1.4 Our contributions
  • We introduce the notion of distribution locality in Definition 2 to capture when weak learning is efficiently achievable by Transformers. The distribution locality measures the least number of tokens required in addition to the token histogram to correlate nontrivially with the target; it is also related in Lemma 6 to correlations with $NC^{0}$ circuits, showing the contrast between learnability and expressivity governed by $TC^{1}$ [16]. It is an explicit measure that applies to a data distribution without requiring an orbit argument to infer a class of distributions [12, 13], giving a tight proxy for models like Transformers. We provide the following results based on the distribution locality:
@@
    • A general conjecture (Conjecture 1), backed by experimental results, that claims efficient weak learning is achievable by a regular Transformer if and only if the distribution locality is constant.
@@
    • A theorem (Theorem 1) that proves the negative side of the above conjecture, the locality barrier, in the instance of a variant of the cycle task under certain technical assumptions. (The cycle task is also put forward in the paper as a simple benchmark to test the global reasoning capabilities of models.)
  • We then switch to the use of ‘scratchpads’ to help with the locality barrier:
@@
    • Agnostic scratchpad: we extend Theorem 1 to cases where a polynomial-size scratchpad is used by the Transformer, without any supervision of the scratchpad. I.e., the scratchpad gives additional memory space for the Transformer to compute intermediate steps. This shows that efficient weak learning is still not possible with such an agnostic scratchpad if the locality is non-constant. An educated guess about what to learn in the scratchpad based on some target knowledge is thus required.
@@
    • Educated scratchpad: we generalize the measure of locality to the ‘autoregressive locality’ to quantify when an educated scratchpad is able to break the locality of a task with subtasks of lower locality. We give experimental results showing that educated scratchpads with constant autoregressive locality allow Transformers to efficiently learn tasks that may originally have high locality. This gives a way to measure how useful a scratchpad can be to break a target into easier sub-targets.
@@
    • We introduce the notion of inductive scratchpad, a type of educated scratchpad that exploits ‘induction’ compared to a fully educated scratchpad. We show that when the target admits an inductive decomposition, such as for the cycle, arithmetic, or parity tasks, the inductive scratchpad both breaks the locality and improves the OOD generalization in contrast to fully educated scratchpads. This gives significant length generalization on additions (from 10 to 20 or from 4 to 26 depending on the method) and parities (from 30 to 50-55). For instance, using different methods, [17] can length generalize from 10 to 13 digits for additions, and [11] can get roughly 10 extra bits for parities with moderate accuracy.
2 Results on the local reasoning barrier
@@
2.1 Defining locality and auto-regressive locality
We define now the notion of distribution locality, which in turn will quantify the notion of locality (or local reasoning) barrier.
@@ 2.1 Defining locality and auto-regressive locality
(Distribution locality) For (a sequence of) distributions $D$ on $\mathcal{A}^{n}\times\mathcal{A}$, where $\mathcal{A}$ is a finite alphabet set of poly$(n)$-cardinality, define $\mathrm{Loc}(D)$ as the smallest number of variables $k\in[n]$ for which there exists $S$, $|S|=k$ such that
@@
The locality $\mathrm{loc}(D)$ of a data distribution measures the least number of input tokens to attend to in order to correlate non-trivially with the label when also given the histogram of the tokens. The specific choice of the mutual information in the definition is not crucial, but one must use a proper measure of dependency (i.e., not just linear correlations), and the mutual information can have convenient chain rule properties. The definition can be related to correlations with $NC^{0}$ circuits (besides for the histogram requirement, see Lemma 6) and also to low-degree polynomial testing, except that it is more general than the latter as it applies to arbitrary token space (without requiring polynomial definitions). Finally, we require the locality to achieve an inverse-polynomial mutual information, the weakest form relevant to weak learning with an inverse-polynomial edge, but one may naturally define the stronger notion with a mutual information of $\Omega(1)$.
We now define the locality in the autoregressive setting.
@@
(Distribution locality in autoregressive setting) For $D$ on $\mathcal{A}^{n}\times\mathcal{A}^{m}$, define $\mathrm{Loc}(D)$ as the smallest integer $k$ for which there exists $S_{1},\ldots,S_{m}$, $|S_{t}|\leq k$ for all $t\in[m]$, such that
@@
In the auto-regressive setting, the locality is mostly relevant when weak learning gives strong learning, in order to let the scratchpad learn each step.
As we will see in the next section, the locality is put forward as a tight proxy to understand efficient weak learning of regular Transformers for arbitrary data distributions. We first present the operational advantages of the definition, going back to the running example of the cycle task.
Attributes of $\mathrm{Loc}$ and some examples.
The locality has the attributes that of being (i) a fairly explicit measure, (ii) applicable to any data distribution on tokens without having to infer a distribution class from the model invariances to estimate the distribution complexity, (iii) not limited to i.i.d. inputs but any input distribution, (iv) relevant to current models of interest such as Transformers.
In particular, back to the cycle task, we have that any set of $n-1$ edges have the same distribution in Class 1 or 2, therefore the locality is at least $n$:
@@
We have Loc(Cycle $\mathrm{Loc}(\text{Cycle task})\geq n$.
As discussed in the next section, this explains why the cycle task is hard to learn. In contrast, the example at the beginning of Section 1.2 has a much lower locality, as being connected correlates to query nodes having large enough degrees, and thus it can be expected for the model to learn with non-trivial accuracy (e.g., using degree shortcuts).
2.2 Transformers require low locality: formal results
@@
    T-regular with $s$-scratchpad if we have a constant sized token alphabet $\Sigma$ and a T-regular Transformer that takes an $m$-sequence in $\Sigma$ and outputs a probability mass function on $\Sigma$ (with well-behaved  [^3] I.e. there exists $c$ such that for any two vectors $x$ and $x^{\prime}$ the total variation distance between the softmax’s probability distributions on these vectors is at most $c||x-x^{\prime}||_{2}$. Note that this does still allow a softmax which returns a specific value with probability $1$ on some vectors. softmax). To compute the value of this net and scratchpad on an input $X\in\Sigma^{n}$ we set $X_{n}=X$ and then for each $m\in\{n,\ldots,n+s-1\}$ draw $x_{m+1}$ from the probability distribution represented by the net’s value on $X_{m}$ and set $X_{m+1}=X_{m}\circ x_{m+1}$ (concatenation). Then, we consider $x_{n+s}$ to be the overall output of the net. [^4] Normally when using a scratchpad we would use causal masking to make it so that we do not need to recalculate previous entries of lists in the Transformer whenever a new entry is added to the scratchpad. We use this one for simplicity, but we still expect all conjectures involving a scratchpad to hold when we apply causal masking to the scratchpad, and Theorem 2 and its proof are unchanged.
@@
In this paper we focus on learnability via descent algorithms, but for the simpler question of expressivity, one wants to know whether there is any choice of parameters for which a Transformer can compute a target function (with limits to how precisely it can record values).
@@
We now state the general conjecture putting forward the locality barrier for learning.
@@
A distribution $P_{X,Y}$ with $P_{X}$ well-behaved  [^6] I.e. $P_{X}$ over $\mathcal{A}^{n}$ such that $|\mathcal{A}|=O(n^{c})$ such that if $X\sim P_{X}$ then there is no value that $X$ takes with probability $O(n^{-c_{1}})$ for any $c_{1}>0$, and every eigenvalue of $X$’s covariance matrix of the indicator functions for the values of $X$’s entries has absolute value $O(n^{-c_{2}}\sum_{i=1}^{n}\mathrm{Var}(X_{i}))$ for some $c_{2}>0$. In other words, the first requirement means that there is no value of $X$ that is frequent enough to weakly learn the function simply by memorizing its value on that input. The second rules out the possibility of an input distribution with a few important components which are easy to notice and determine the output, such as a scenario where there are $\log(n)$ blocks of $n/(2\log(n))$ bits such that all of the bits in a block are always the same and the output depends only on the values of the blocks. Other than that, many distributions of interest are well-behaved, such as the iid measure as in [12, 25, 24, 28, 26, 27], or measures with dependencies as for the cycle task considered here. is efficiently weakly learnable (inverse polynomial edge) by a T-regular Transformer only if $P_{X,Y}$ has constant locality. Conversely, if $|\mathcal{A}|$ is bounded and $P_{X,Y}$ has constant locality then $P_{X,Y}$ is efficiently weakly learnable by a T-regular Transformer.
@@
(1) An important property of the learning model for the above conjecture is that the probability distribution of the function computed by the model is invariant under permutations of the inputs, and if it is trained in a reasonable way on samples drawn from a probability distribution drawn from a class that is symmetric under permutations of the inputs, its probability distribution will retain its symmetry under permutations of the inputs. For MLPs, we expect most of the results in this paper to apply, with the modification of the locality not having access to the empirical measure of $X$, since one has additional symmetry obtained by swapping tokens. (2) If we were to use causal masking or a specific choice of positional embeddings that would essentially make it easier for the Transformer to focus on specific subsets of the inputs, and if such subsets are relevant, which may be engineered using some prior knowledge of the target, one could potentially learn functions with higher locality. For instance, we would expect to be unable to learn a function that computes the parity of some subset of $\log(n)$ input bits. However, if we had a positional embedding that gave one value for all of the active bits and mapped all of the other bits to 00, then the Transformer would probably be able to learn the function in question. Likewise, causal masking makes it so that the first few elements of any list the Transformer computes depend only on the first few input symbols, which makes it easier to learn functions that would rely on those symbols.
@@
2.3 Agnostic scratchpads cannot break the locality

### v2

# How Far Can Transformers Reason? The Globality Barrier and Inductive Scratchpad

@@ Abstract
Can Transformers predict new syllogisms by composing established ones? More generally, what type of targets can be learned by such models from scratch? Recent works show that Transformers can be Turing-complete in terms of expressivity, but this does not address the learnability objective. This paper puts forward the notion of globality degree of a target distribution to capture when weak learning is efficiently achievable by regular Transformers, where the latter measures the least number of tokens required in addition to the tokens histogram to correlate nontrivially with the target. As shown experimentally and theoretically under additional assumptions, distributions with high globality cannot be learned efficiently. In particular, syllogisms cannot be composed on long chains. Furthermore, we show that (i) an agnostic scratchpad cannot help to break the globality barrier, (ii) an educated scratchpad can help if it breaks the globality at each step, however not all such scratchpads can generalize to out-of-distribution (OOD) samples (iii) a notion of ‘inductive scratchpad’, that composes the prior information more efficiently, can both break the globality barrier and improve the OOD generalization. In particular, some inductive scratchpads can achieve length generalizations of up to 6×6$\times$6 × for some arithmetic tasks depending on the input formatting.
@@ 1.2 Hardness of long compositions
Consider the previous syllogism composition task where implications are drawn on a graph with 24 edges drawn randomly over 24 vertices. Picking vertices at distances 1 to 4 for the connected case and picking disconnected vertices uniformly at random lets a Transformer achieve a test accuracy of more than 80% after about 2000 iterations. However, does this mean that the model has learned to compose syllogisms, or has it found shortcuts, e.g., based on node degrees, to guess the implications often enough? In Appendix B.1, we provide empirical evidence supporting the latter. Motivated by this issue and to preclude spurious correlations, we consider the following distribution.
@@
(b) Nb. of iterations (y-axis) with $512$ fresh samples to learn ($\geq 95\%$ acc.) the cycle task of size $n$ (x-axis) using GPT2-style models with $10M,25M,85M$ parameters.
@@
A challenge for the cycle task is that there is no clear ‘low-complexity pattern’ in the input representation that indicates whether there are 1 or 2 cycles. No simple statistics based on degrees, edge counts, or finite motif counts that can tell if the vertices are connected or not. One has to consider at least $n$ edges in order to get any correlation with the presence of a path. In other words, the task requires a ‘global reasoning’ involving a ‘large’ number of input tokens and this seems to be hard for Transformers.
@@ 1.3 Hardness of global reasoning
As discussed previously, the cycle task appears to be challenging for Transformers as it requires some global reasoning. Other tasks such as subset parities exhibit the same challenge. However the latter can be proved to be not efficiently learnable by various regular neural networks and noisy gradient descent, as one can get explicitly a class of functions (through orbit arguments [12, 13]) that has large statistical dimension [14] or low cross-predictability [12, 15] (see Appendix A.2). For the cycle task, we have a single distribution, and it is unclear how to use the invariances of Transformers to get arguments as in [12, 13], as the input distribution is not invariant under the symmetries of the model. We thus would like to develop a more general complexity measure that unifies why such tasks are hard for Transformer-like models and that formalizes the notion of ‘global reasoning barrier’ when models are trained from scratch. We also would like to understand how the scratchpad methodologies that have proved helpful in various settings (see Section 3) can help here. This raises the questions:
(1) How can we formalize the ‘global reasoning barrier’ in general terms?
(2) Can we break the ‘global reasoning barrier’ with scratchpad methodologies?
@@ 1.4 Our contributions
  • We introduce the notion of globality degree in Definition 2 to capture when weak learning is efficiently achievable by Transformers. The globality degree measures the least number of tokens required in addition to the token histogram to correlate nontrivially with the target; it is also related in Lemma 6 to correlations with $NC^{0}$ circuits, showing the contrast between learnability and expressivity controlled by $TC^{0}/TC^{1}$ with constant/logarithmic depth [16]. It is an explicit measure that applies to a data distribution without requiring an orbit argument to infer a class of distributions [12, 13], giving a tight proxy for models like Transformers. We provide the following results based on the globality degree:
@@
    • A general conjecture (Conjecture 1), backed by experimental results, that claims efficient weak learning is achievable by a regular Transformer if and only if the globality degree is constant.
@@
    • A theorem (Theorem 1) that proves the negative side of the above conjecture, the globality barrier, in the instance of a variant of the cycle task under certain technical assumptions. (The cycle task is also put forward in the paper as a simple benchmark to test the global reasoning capabilities of models.)
  • We then switch to the use of ‘scratchpads’ to help with the globality barrier:
@@
    • Agnostic scratchpad: we extend Theorem 1 to cases where a polynomial-size scratchpad is used by the Transformer, without any supervision of the scratchpad. I.e., the scratchpad gives additional memory space for the Transformer to compute intermediate steps. This shows that efficient weak learning is still not possible with such an agnostic scratchpad if the globality is non-constant. An educated guess about what to learn in the scratchpad based on some target knowledge is thus required.
@@
    • Educated scratchpad: we generalize the measure of globality to the ‘autoregressive globality’ to quantify when an educated scratchpad is able to break the globality of a task with subtasks of lower globality. We give experimental results showing that educated scratchpads with constant autoregressive globality allow Transformers to efficiently learn tasks that may originally have high globality. This gives a way to measure how useful a scratchpad can be to break a target into easier sub-targets.
@@
    • We introduce the notion of inductive scratchpad, a type of educated scratchpad that exploits ‘induction’ compared to a fully educated scratchpad and thus composes more efficiently the prior state information. We show that when the target admits an inductive decomposition, such as for the cycle, arithmetic, or parity tasks, the inductive scratchpad both breaks the globality and improves the OOD generalization in contrast to fully educated scratchpads. This gives significant length generalization on additions (from 10 to 20 or from 4 to 26 depending on the method) and parities (from 30 to 50-55). For instance, using different methods, [17] can length generalize from 10 to 13 digits for additions, and [11] can get roughly 10 extra bits for parities with moderate accuracy.
2 Results on the global reasoning barrier
@@
2.1 Defining globality and auto-regressive globality
We define now the notion of globality degree, which in turn will quantify the notion of globality (or global reasoning) barrier.
@@ 2.1 Defining globality and auto-regressive globality
(Globality degree) For (a sequence of) distributions $D$ on $\mathcal{A}^{n}\times\mathcal{A}$, where $\mathcal{A}$ is a finite alphabet set of poly$(n)$-cardinality, define the globality degree of $D$, $\mathrm{glob}(D)$, as the smallest number of variables $k\in[n]$ for which there exists $S$, $|S|=k$ such that
@@
The globality degree, or simply globality, of a data distribution measures the least number of input tokens to attend to in order to correlate non-trivially with the label when also given the histogram of the tokens. The specific choice of the mutual information in the definition is not crucial, but one must use a proper measure of dependency (i.e., not just linear correlations), and the mutual information can have convenient chain rule properties. The definition can be related to correlations with $NC^{0}$ circuits (besides for the histogram requirement, see Lemma 6) and also to low-degree polynomial testing, except that it is more general than the latter as it applies to arbitrary token space (without requiring polynomial definitions). Finally, we require the globality to achieve an inverse-polynomial mutual information, the weakest form relevant to weak learning with an inverse-polynomial edge, but one may naturally define the stronger notion with a mutual information of $\Omega(1)$.
We now define the globality in the autoregressive setting.
@@
(Globality degree in autoregressive setting) For $D$ on $\mathcal{A}^{n}\times\mathcal{A}^{m}$, define $\mathrm{glob}(D)$ as the smallest integer $k$ for which there exist sets of indices $S_{1},\ldots,S_{m}$, $|S_{t}|\leq k$ for all $t\in[m]$, such that
@@
In the auto-regressive setting, the globality is mostly relevant when weak learning gives strong learning, in order to let the scratchpad learn each step.
As we will see in the next section, the globality degree is put forward as a tight proxy to understand efficient weak learning of regular Transformers for arbitrary data distributions. We first present the operational advantages of the definition, going back to the running example of the cycle task.
Attributes of $\mathrm{glob}$ and some examples.
The globality has the attributes of being (i) a fairly explicit measure, (ii) applicable to any data distribution on tokens without having to infer a distribution class from the model invariances to estimate the distribution complexity, (iii) not limited to i.i.d. inputs but any input distribution, (iv) relevant to current models of interest such as Transformers.
In particular, back to the cycle task, we have that any set of $n-1$ edges have the same distribution in Class 1 or 2, therefore the globality is at least $n$:
@@
We have glob(Cycle $\mathrm{glob}(\text{Cycle task}(n))\geq n$.
As discussed in the next section, this explains why the cycle task is hard to learn. In contrast, the example at the beginning of Section 1.2 has a much lower globality, as being connected correlates to query nodes having large enough degrees, and thus it can be expected for the model to learn with non-trivial accuracy (e.g., by using degree shortcuts).
2.2 Transformers require low globality: formal results
@@
    T-regular with $s$-scratchpad if we have a constant sized token alphabet $\Sigma$ and a T-regular Transformer that takes an $m$-sequence in $\Sigma$ and outputs a probability mass function on $\Sigma$ (with well-behaved [^3] I.e. there exists $c$ such that for any two vectors $u$ and $u^{\prime}$ the total variation distance between the softmax’s probability distributions on these vectors is at most $c||u-u^{\prime}||_{2}$. Note that this does still allow a softmax which returns a specific value with probability $1$ on some vectors. softmax). To compute the value of this net and scratchpad on an input $X\in\Sigma^{n}$ we set $X_{n}=X$ and then for each $m\in\{n,\ldots,n+s-1\}$ draw $x_{m+1}$ from the probability distribution represented by the net’s value on $X_{m}$ and set $X_{m+1}=X_{m}\circ x_{m+1}$ (concatenation). Then, we consider $x_{n+s}$ to be the overall output of the net. [^4] Normally when using a scratchpad we would use causal masking to make it so that we do not need to recalculate previous entries of lists in the Transformer whenever a new entry is added to the scratchpad. We use this one for simplicity, but we still expect all conjectures involving a scratchpad to hold when we apply causal masking to the scratchpad, and Theorem 2 and its proof are unchanged.
@@
In this paper, we focus on learnability via descent algorithms, but for the simpler question of expressivity, one wants to know whether there is any choice of parameters for which a Transformer can compute a target function (with limits to how precisely it can record values).
@@
We now state the general conjecture putting forward the globality barrier for learning.
@@
A distribution $P_{X,Y}$ with well-behaved [^6] I.e. $P_{X}$ over $\mathcal{A}^{n}$ such that $|\mathcal{A}|=O(n^{c})$ such that if $X\sim P_{X}$ then there is no value that $X$ takes with probability $O(n^{-c_{1}})$ for any $c_{1}>0$, and every eigenvalue of $X$’s covariance matrix of the indicator functions for the values of $X$’s entries has absolute value $O(n^{-c_{2}}\sum_{i=1}^{n}\mathrm{Var}(X_{i}))$ for some $c_{2}>0$. In other words, the first requirement means that there is no value of $X$ that is frequent enough to weakly learn the function simply by memorizing its value on that input. The second rules out the possibility of an input distribution with a few important components that are easy to notice and determine the output, such as a scenario where there are $\log(n)$ blocks of $n/(2\log(n))$ bits such that all of the bits in a block are always the same and the output depends only on the values of the blocks. Other than that, many distributions of interest are well-behaved, such as the i.i.d. measure as in [12, 25, 24, 28, 26, 27], or measures with dependencies as for the cycle task considered here. $P_{X}$ is efficiently weakly [^7] I.e., with an accuracy that improves on the trivial accuracy by at least $n^{-c}$ for a constant $c$. learnable by a T-regular Transformer if and only if [^8] More specifically, for the converse statement assume that $|\mathcal{A}|$ is constant. $P_{X,Y}$ has constant globality.
@@
(1) An important property of the learning model for the above conjecture is that the probability distribution of the function computed by the model is invariant under permutations of the inputs, and if it is trained in a reasonable way on samples drawn from a probability distribution drawn from a class that is symmetric under permutations of the inputs, its probability distribution will retain its symmetry under permutations of the inputs. For MLPs, we expect most of the results in this paper to apply, with the modification of the globality not having access to the empirical measure of $X$, since one has additional symmetry obtained by exchanging tokens. (2) If we were to use causal masking or a specific choice of positional embeddings that would essentially make it easier for the Transformer to focus on specific subsets of the inputs, and if such subsets are relevant, which may be engineered using some prior knowledge of the target, one could potentially learn functions with higher globality. For instance, we would expect to be unable to learn a function that computes the parity of some subset of $\log(n)$ input bits. However, if we had a positional embedding that gave one value for all of the active bits and mapped all of the other bits to 00, then the Transformer would probably be able to learn the function in question. Likewise, causal masking makes it so that the first few elements of any list the Transformer computes depend only on the first few input symbols, which makes it easier to learn functions that would rely on those symbols.
@@
2.3 Agnostic scratchpads cannot break the globality